### v1

# Benchmarking Cross-Domain Audio-Visual Deception Detection

Abstract
Automated deception detection is crucial for assisting humans in accurately assessing truthfulness and identifying deceptive behavior. Conventional contact-based techniques, like polygraph devices, rely on physiological signals to determine the authenticity of an individual’s statements. Nevertheless, recent developments in automated deception detection have demonstrated that multimodal features derived from both audio and video modalities may outperform human observers on publicly available datasets. Despite these positive findings, the generalizability of existing audio-visual deception detection approaches across different scenarios remains largely unexplored. To close this gap, we present the first cross-domain audio-visual deception detection benchmark, that enables us to assess how well these methods generalize for use in real-world scenarios. We used widely adopted audio and visual features and different architectures for benchmarking, comparing single-to-single and multi-to-single domain generalization performance. To further exploit the impacts using data from multiple source domains for training, we investigate three types of domain sampling strategies, including domain-simultaneous, domain-alternating, and domain-by-domain for multi-to-single domain generalization evaluation. Furthermore, we proposed the Attention-Mixer fusion method to improve performance, and we believe that this new cross-domain benchmark will facilitate future research in audio-visual deception detection. Protocols and source code are available at https://github.com/Redaimao/cross_domain_DD.
Index Terms:
audio-visual, multimodal deception detection, cross-domain, generalization.
1 Introduction
Figure 1: Typical samples from different publicly deception detection datasets: Real Life Trials [1], Bag of Lies [2], MU3D [3], and Box of Lies [4]. The samples in each row are from different datasets while those in each column are with different modalities (visual vs. audio) and ground truth labels (i.e., truthful vs. deceptive). It can be seen that serious domain shifts (e.g., resolution/illumination/pose in visual faces and pitch/loudness/noise in audio) occur among these datasets.
Audio-visual deception detection involves utilizing AI techniques and algorithms to automatically detect deceptive behavior in speech and facial movements [5, 6, 7, 8]. Deception detection has a significant impact on various real-world applications such as law enforcement [9], healthcare [10], and business [11]. It has the potential to prevent fraud, improve security measures, and enhance trust and confidence. A reliable deception detection tool can support more accurate decision-makings.
Traditional deception detection is often a contact-based method. It assesses whether someone is telling the truth or not by monitoring physiological responses like skin conductance and heart rate [12]. Experts’ Behavioral observation and analysis are another technique that evaluates changes in a person’s body language, speech patterns, and eye movements [13, 14]. However, such an assessment can be time-consuming and require significant expertise to perform accurately.
Recently, the development of automated deception detection systems using AI and machine learning techniques has gained significant attention as the existing methods have limitations in terms of reliability, accuracy, and scalability. Various multimodal datasets have been introduced, including real-life trials from court scenes [1], lab-based setups [2, 3], and game show scenarios [4]. These datasets provide a wide variety of deceptive samples from different domains, enabling researchers to examine the effectiveness of AI models on deception detection. Based on these datasets, progress has been made in deception detection techniques within specific domains [7, 6, 15]. Recent studies have utilized rich visual and audio features [16, 17, 18], such as Mel Spectrogram, emotional states, and facial action units, to enhance the performance of deception detection tasks.
However, there remains a substantial research gap that needs to be addressed. Specifically, fewer studies have explored the cross-domain issue, despite the presence of significant domain shifts in public deception detection datasets. As shown in Figure 1, domain shifts are observed in both audio and visual modalities from publicly available datasets. The generalizability of the models is critical for practical applications. Therefore, such domain shifts need to be investigated in order to develop deception detection models that can be generalized across different contexts. Additionally, effective methods must be proposed to alleviate the domain shift issue by fusing both audio and visual features in a meaningful way. Addressing these issues can benefit automated deception detection systems in improving generalizability in real-world applications.
To address the issue of cross-domain deception detection, we introduce a new benchmark that evaluates the generalization capacity of AI models using audio and visual features over publicly available datasets. Our benchmarking approach utilizes widely adopted audio and visual features, and we compare the single-to-single domain performance and multi-to-single domain generalization using different architectures. Specifically, for the multi-to-single setting, three domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, are implemented to conduct cross-domain testing. To further enhance performance, we propose an Attention-Mixer fusion method based on MLP-Mixer [19]. This benchmarking framework serves as an important tool for evaluating the effectiveness of audio-visual deception detection models in diverse contexts, which will help improve the capabilities of automated deception detection systems in real-world settings. Additionally, we hope our work will inspire further research on multimodal models that address domain shift issues. In summary, our main contributions include:
  • Introducing a new benchmark for evaluating the generalization capacity of AI models using audio and visual features across different domains.
  • Comparing the single-to-single domain and multi-to-single domain generalization using different architectures.
  • Providing three domain sampling strategies i.e., domain simultaneous, domain alternating, and domain-by-domain, to conduct multi-to-single cross-domain testing.
  • Proposing the Attention-Mixer fusion method to enhance performance.
In the rest of the paper, Sec. 2 provides a review of related psychological studies on cues to deception and multimodal deception detection works. Sec. 3 introduces our benchmarking approach and fusion method. Sec. 4 provides the cross-domain benchmark results and fusion results. Finally, conclusions and future works are given in Sec. 5.
2 Related work
2.1 Cues to Deception
The research on using behavioral cues for deception has gradually become active over the past few decades. Psychological researchers have published a large number of works on the analysis of cues to deception [20, 21, 22]. Among the studied behavioral cues, verbal and nonverbal cues were preferred as humans may behave differently between lying and telling the truth. DePaulo et al. [23] studied and reported experimental results on 158 cues to deception. They revealed that, in general, people who tell lies are less forthcoming and less convincing than those who tell the truth. Liars usually talk about fewer details and make fewer spontaneous corrections. They also sound less involved but more vocally tense. Through the study, the researchers statistically found that liars often press their lips, repeat words, raise their chins, and show less genuine smiles. The results show that some behavioral cues do potentially appear in deception and are even more pronounced when liars are more motivated to cheat.
Levine et al. [24] reviewed the status quo and provided a new perspective on the theories of deception. They pointed out that lying usually happens when problematic information is involved. It is critical to understand the verbal content in the context. Vrij et al. [25] realized that interviewers play a vital role in eliciting and enhancing cues to deceit. The authors proposed the “interviewing to detect deception” technique to open a new path in the deception detection research field. They argued that different psychological states can be exploited by adopting appropriate interview techniques of liars and truth-tellers. Hirschberg et al. [26] proposed a method to distinguish deceptive from non-deceptive speeches by a large corpus. They also conducted experiments using acoustic, lexical, and speaker-dependent features, which showed improved performance by combining multiple feature sets. Warren et al. [27] conducted experiments to investigate the relationship between affective facial expressions to deception. The results indicated that leaked emotions with the incongruous intended message can provide useful cues to deception, which supported the nonverbal leakage theory [28, 29].
Figure 2: Main network and method. (a) Model architecture. Visual modality includes face and behavior inputs. Audio modality includes Mel Spectrogram input. The features obtained by the respective encoders. The fusion methods include score fusion and feature fusion. (b) Domain sampling strategies. Domain-simultaneous: each batch consists of samples from multiple sources. Domain-alternating: each batch is alternatively sampled from multiple sources. Domain-by-domain: the batches are sampled from one source and then from another.
2.2 Multimodal Deception Detection
Recent works for deception detection usually use verbal and non-verbal features and propose effective fusion methods [30, 15, 16, 17]. For example, some works utilized facial features from RGB images to perform deception detection [15, 17, 5]. To capture facial movements, facial action units (AUs) were utilized. Other features, such as facial expression, were also adopted [6, 30]. Besides visual features, many works incorporated audio features to boost performance [7, 6, 15]. For example, Wu et al. [6] used MFCC (Mel-frequency Cepstral Coefficients) features and Karimi et al. [15] used raw audio. Most of the recent works mentioned above have considered multimodal fusion approaches that extract visual, audio, and text information to boost performance. In addtion to visual, audio and text, Karnati et al. [16] exploited physiological signals, i.e., EEG representations for deception detection.
To better fuse the multimodal features, different fusion methods were proposed, which can be broadly categorized into feature-level fusion and decision-level fusion. Specifically, feature-level fusion focused on producing better multimodal embeddings and used the linear layers to extract crossmodal dynamics [15, 17, 5, 6, 7]. Rather than that, decision-level fusion aimed to fuse multimodal dynamics at a late stage, to reduce computational complexity and learn good marginal representations [16, 7].
However, previous works on multimodal deception detection did not consider cross-domain issues that occur from one domain to another, which is the focus of this work.
3 Methodology
The mainstream architecture for audio-visual deception detection usually includes encoders for unimodal feature extraction and/or a fusion module. We follow the widely adopted architecture to build the benchmark on cross-domain audio-visual deception detection in this work. As shown in Fig. 2, audio and visual features are extracted from audio and visual encoders. The fusion module is performed based on audio and visual features. The fused feature is input to the classifier for classification. We build the benchmark for cross-domain generalization performance based on such network architecture with different encoders. We conducted single-to-single and multi-to-single evaluations where three domain sampling strategies included, i.e., domain-simultaneous, domain-alternating, and domain-by-domain.
3.1 Audio and Visual Feature Learning
To establish a benchmark for cross-domain audio-visual deception detection, we utilize widely adopted audio and visual features along with their respective encoders. Our approach treats audio and visual features as equally important, extracting different types of features simultaneously. As depicted in Fig. 1, this network structure offers several advantages: (1) flexibility in network selection: different audio or visual encoders can be effortlessly incorporated and compared in a fair manner, (2) adaptability: the addition or removal of specific modules and/or losses is straightforward. For instance, a fusion module can be inserted before classifiers, and (3) easy performance benchmarking: the system facilitates evaluating performance in various settings, such as score-level fusion and feature-level fusion. In this work, we focus on audio and visual modalities for deception detection. In particular, two kinds of visual features are extracted, i.e., face features from RGB face images and behavior features consist of AUs, affects, etc.
As shown in Fig. 1, given a detected RGB face image as input $X_{f}$, the deep features $F_{f}$ could be extracted via face encoder networks $\mathcal{E}_{f}$ (e.g., ResNet18 [31]). Similarly, behavior inputs such as the AU and/or affect features $X_{b}$ are encoded by OpenFace [32] or affect model (e.g., EmotionNet [33]) $\mathcal{E}_{b}$ to output behavior features $F_{b}$. Note that we regard both face frames and behavior features as the visual modality but differentiate them in this work as they have different types of information and representations. Given audio input $X_{a}$ (either Mel Spectrogram [34] or waveforms), audio features $F_{a}$ are extracted through audio encoder $\mathcal{E}_{a}$. The corresponding classification heads for face frames ($\mathcal{H}_{f}$), behavior features ($\mathcal{H}_{b}$), and audio features ($\mathcal{H}_{a}$) output the prediction logits $\hat{Y}_{f}$, $\hat{Y}_{b}$, and $\hat{Y}_{a}$, respectively. The fusion head $\mathcal{G}$ takes $F_{f}$, $F_{b}$, and $F_{a}$ as input. $\mathcal{G}$ is determined by the actual fusion method, e.g., liner layer, transformer layers, MLP, etc. The output logit of $\mathcal{G}$ is denoted by $\hat{Y}_{g}$. Therefore, the audio and visual learning process can be denoted as follows:
$\small\begin{split}F_{f}&=\mathcal{E}_{f}(X_{f}),\hat{Y}_{f}=\mathcal{H}_{f}(F_{f}),\\ F_{b}&=\mathcal{E}_{b}(X_{b}),\hat{Y}_{b}=\mathcal{H}_{b}(F_{b}),\\ F_{a}&=\mathcal{E}_{a}(X_{a}),\hat{Y}_{a}=\mathcal{H}_{a}(F_{a}),\\ \hat{Y}_{g}&=\mathcal{G}(F_{f},F_{b},F_{a}).\end{split}$ (1)
Loss Function. For deception detection ground truth $Y$, where $Y=0$ for truthful and $Y=1$ for deception, the binary cross-entropy loss (BCE) is adopted. The loss for each sample with a certain modality or fused prediction can be denoted as
$\mathcal{L}_{m}=-(Ylog(\hat{Y}_{m})+(1-Y)log(1-\hat{Y}_{m})),$ (2)
where $m\in\{f,b,a,g\}$, $\hat{Y}_{m}$ is the corresponding prediction logits. In other words, the BCE loss is calculated separately for each type of modality and/or its fused feature depending on whether a sample has any face frames, visual inputs, or audio inputs. The overall loss function can be described as follows:
$\mathcal{L}=\frac{1}{N}\sum_{i=i}^{N}\left(\sum_{m={f,b,a}}\mathcal{L}_{m,i}+\lambda\mathcal{L}_{g,i}\right),$ (3)
where $N$ is the number of data samples and $\lambda$ is a trade-off parameter between modality loss and fusion loss. $\lambda$ is set to 0.5 in our experiments.
3.2 Cross-domain Generalization
We benchmark the cross-domain generalization on the deception detection task. First, we introduce the notations and definitions in this section. A domain is composed of data that are sampled from a distribution (dataset), which can be denoted as $\mathcal{S}=\{(X;Y)_{i}\}_{i=1}^{N}\sim P_{S}$, where $X=(X_{f},X_{b},X_{a})$, $X_{f},X_{b},X_{a}$ represent samples of face frames, behavior, and audio modalities, respectively. $Y$ denotes the label, and $P_{S}$ denotes the joint distribution of the input samples and the output label. In this paper, for simplicity, we follow similar definitions in [35, 36] to treat each dataset as an individual domain due to their obvious distribution gaps, but more fine-grained intra-domain factors would be explored in future work. For domain generalization, $M$ source domains (training datasets) are given, i.e., $\mathcal{S}_{train}=\{S_{j}|j=1,\cdots,M\}$, where $\mathcal{S}_{j}=\{(X;Y)_{i}\}_{i=1}^{N_{j}}\sim P_{S_{j}}$ denotes the $j$-th domain, and $P_{S_{i}}\neq P_{S_{j}}$ for $1\leq i,j\leq M$. $N_{j}$ is the number of total samples in $S_{j}$. The goal of domain generalization is to learn the predictive function $h$ in $M$ source domains to achieve minimum error on an unseen test domain $\mathcal{S}_{test}\sim P_{S_{test}}$, and $P_{S_{test}}\neq P_{S_{j}}$ for $1\leq j\leq M$:
$min~{}\mathbb{E}_{(X;Y)\in\mathcal{S}_{test}}\left[\mathcal{L}(h(X),Y)\right],$ (4)
where $X=(X_{f},X_{b},X_{a})$, $Y$ is the label, $\mathcal{L}$ is the loss function, and $\mathbb{E}$ is the expectation.
When $M=1$, it is a Single-to-single Cross-domain Generalization task, where the modal is trained on one training dataset and tested on another dataset. When $M\geqslant 2$, we propose three strategies to learn from multiple domains for the Multi-to-single Cross-domain Generalization. Let $B$ denote one batch of training data with a size of $N_{B}$. Given multiple training domains $\mathcal{S}_{train}=\{S_{j}|j=1,\cdots,M\}$, $B$ is a set of training data sampled from $\mathcal{S}_{train}$.
Domain-Simultaneous means to train multiple domains in parallel within each batch of data. In domain simultaneous training, the $k-$th batch of training data is a group of samples from different domains, i.e., $B^{k}=(b_{S_{1}}^{k},\cdots,b_{S_{M}}^{k})$, $k\in[1,\cdots K]$, where $b_{S_{j}}^{k}$ is the batch samples from domain $S_{j}$ for $j=1,\cdots,M$, $K$ is the number of batches during training. The total number of $b_{S_{j}}^{k}$ is $N_{B}$. As shown in Fig. 2 (b), each training batch contains smaller batch samples from all the source domains during training. Models are trained to learn from different domains simultaneously by feeding the mixed batch data.
[…]
3.3 Attention-Mixer Fusion
Besides investigating cross-domain sampling strategies, inspired by [19], we propose Attention-Mixer Fusion to enhance the performance by fusing audio-visual modalities, where the attention mixer layer takes multimodal features as input to produce fused features. In particular, an attention mixer layer is composed of unimodal MLP layers, self-attention layers [37], and crossmodal MLP layers. First, for batch size $N_{B}$, the input features from different modalities are concatenated and projected to be a tensor $U\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ by a Liner layer, followed by several attention mixer layers, where $N_{m}$ is the number of input modalities. Specifically, the unimodal MLP layer, the self-attention layer, and the crossmodal MLP layer can be respectively described as
$\small U^{*,*,i}=F_{g}^{*,*,i}+\mathbf{W_{2}}\ \sigma(\mathbf{W_{1}}\ LN(F_{g}^{*,*,i})),\ i=\left[1,D\right],$ (5)
$\footnotesize U=\left[\left(softmax\left(U{\mathbf{W_{3}}(U\mathbf{W_{4}})^{T}}\over{\sqrt{D}}\right)U\mathbf{W_{5}}\right)_{h}\right]\mathbf{W_{6}},\ h=[1,H],$ (6)
$\small U^{*,j,*}=U^{*,j,*}+\mathbf{W_{8}}\ \sigma(\mathbf{W_{7}}\ LN(U^{*,j,*})),\ j=\left[1,N_{m}\right],$ (7)
where $LN(\cdot)$ denotes the Layer Normalization, $\mathbf{W}_{1-8}$ are trainable weights, $H$ is the number of heads in multihead self-attention, and * denotes all the entries in that dimension. Several attention mixer layers are stacked as a deep block, which is set as a hyperparameter in practice. We set it to 6 in our experiment. Finally, $U\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ is reduced to $U\in\mathbb{R}^{N_{B}\times N_{m}\times 1}$ by obtaining the mean value on the feature dimension. In Eq. 5, the unimodal MLP layer is conducted along the feature dimension to learn the dynamics in each unimodal feature. Eq. 6 shows the multi-head self-attention operation on the tensor $U$, which further explores the attention between the unimodal features. In Eq. 7, the crossmodal MLP layer learns the dynamics across the modality dimension from the corresponding feature tokens.
TABLE I: The results of single-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M).
Modality & Inputs Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
Visual (AU) LSTM [38] 48.11 - - 61.21 - - - - - -
Visual (Face frames) ResNet18 52.00 61.39 51.25 50.93 57.43 50.62 57.94 51.69 57.43 54.52
Visual (Face frames) ResNet18+GRU 53.54 63.37 52.81 57.41 59.41 51.56 46.73 52.92 55.45 54.80
Visual (AU+Gaze) MLP 50.77 65.35 56.87 58.88 58.42 50.94 46.73 51.69 53.47 54.79
Visual (Affect) MLP 50.46 58.42 50.31 50.47 51.49 52.19 66.36 51.08 60.40 54.58
Visual (AU+Gaze+Affect) MLP 54.46 59.41 54.37 50.47 57.43 54.69 60.75 51.69 55.45 55.41
Audio (Mel spectrogram) ResNet18 46.77 53.47 52.19 50.47 66.34 50.62 54.21 51.38 55.45 53.43
Audio (Waveform) Wave2Vec 51.08 48.51 50.94 46.73 58.42 50.00 63.55 56.31 56.44 53.55
TABLE II: The results of multi-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M), for different generalization strategies.
Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
Domain-Simultaneous
Visual (Face frames) ResNet18 53.85 49.50 49.50 50.94 44.86 60.40 44.55 50.51
Visual (Face frames) ResNet18+GRU 52.62 54.46 51.49 51.88 53.27 59.41 44.55 52.53
Visual (AU+Gaze) MLP 53.54 47.52 48.51 50.94 52.34 53.47 56.44 51.82
Visual (Affect) MLP 50.15 54.46 55.45 52.19 53.27 57.43 62.38 55.04
Visual (AU+Gaze+Affect) MLP 50.46 52.48 61.39 51.25 51.4 60.4 63.37 55.82
Audio (Mel spectrogram) ResNet18 48.92 45.54 53.47 53.12 43.93 62.38 50.5 51.12
Audio (Waveform) Wave2Vec 52.92 55.45 44.55 51.25 69.16 42.57 46.53 51.78
Domain-Alternating
Visual (Face frames) ResNet18 50.15 45.54 56.44 51.56 50.47 54.46 65.85 53.50
Visual (Face frames) ResNet18+GRU 55.38 52.48 60.40 50.00 50.47 60.40 64.62 56.25
Visual (AU+Gaze) MLP 55.45 47.52 53.47 51.25 54.21 57.43 60.40 54.25
Visual (Affect) MLP 51.08 56.44 61.39 52.19 52.34 58.42 53.47 55.05
Visual (AU+Gaze+Affect) MLP 51.38 58.42 63.37 50.31 53.27 52.48 60.40 55.66
Audio (Mel spectrogram) ResNet18 50.15 60.40 53.47 50.31 58.88 51.49 47.52 53.17
Audio (Waveform) Wave2Vec 52.92 55.45 44.55 50.62 64.49 58.42 48.51 53.57
Domain-by-Domain
Visual (Face frames) ResNet18 52.00 53.47 56.44 50.00 59.81 41.58 55.45 52.68
Visual (Face frames) ResNet18+GRU 54.46 41.58 66.34 50.62 51.40 56.44 60.40 54.46
Visual (AU+Gaze) MLP 51.08 43.56 55.45 53.75 57.01 53.47 54.46 52.68
Visual (Affect) MLP 55.69 57.43 57.43 51.56 52.34 49.50 61.39 55.05
Visual (AU+Gaze+Affect) MLP 50.15 56.44 58.42 50.00 57.94 60.40 63.37 56.67
Audio (Mel spectrogram) ResNet18 52.31 50.50 58.42 49.38 53.27 56.44 59.41 54.24
Audio (Waveform) Wave2Vec 56.00 47.52 44.55 53.12 67.29 57.43 58.42 54.90
4 Experiments
In this part, extensive experiments are conducted to benchmark the cross-domain performances on public deception detection datasets. In the following, we sequentially describe the benchmark datasets & metrics (Sec. 4.1), implementation details (Sec. 4.2), benchmarking results (Sec. 4.3 - 4.4) and fusion performances (Sec. 4.5).
4.1 Databases and Metrics
Datasets. We benchmarked the cross-domain generalization performance based on four publicly available datasets. Real Life Trials [1] dataset is a popular real-world dataset collected from public court trials, which consists of 121 videos including 61 deceptive and 60 truthful video clips. As it is a real-world dataset, the Real Life trial dataset has more noise on both the video and audio. We filtered out some corrupted videos and obtained 108 videos (54 truthful and 54 deceptive) with 58 subjects for our experiments. Bag of Lies [2] is a multimodal dataset collected from well-controlled lab-based scenarios, where video, audio, EEG, and gaze data are collected. It has 35 subjects, 163 truthful and 162 deceptive video clips. The backgrounds for the videos are relatively clean and it is less noisy. MU3D [3] has 320 video clips and 80 subjects that cover different races and genders. It is also a lab-based dataset that uses the personal description paradigm to stimuli real-world cases. Each participant tells a positive truth, a positive lie, a negative truth, and a negative lie. Box of Lies [4] is a deception dataset collected from an online gameshow, which has 25 videos and 26 participants (6 male and 20 female). The full video set contains 29 truthful and 36 deceptive rounds of games. However, the quality of the original Box of Lies dataset is not satisfactory. The visual (the face of the participant) and audio from many clips are not matching due to the frequent changes of viewpoints. To perform a fair comparison, we preprocessed and cleaned the Box of Lies dataset. After preprocessing, 101 video clips were extracted for testing. Some of the typical samples from these datasets are shown in Fig. 1.
Evaluation Metrics. In this work, we followed the widely adopted metric, binary classification accuracy (%), for experimental evaluation. The deceptive clips were labeled as 1 and truthful clips were labeled as 0.
4.2 Implementation Details
Feature Extraction. Several widely-adopted audio and visual features were extracted by different tools. For visual features, OpenFace [32] was used to extract 35-dimensional AUs and 8-dimensional gaze features. Face frames were extracted and aligned by MTCNN [39], where we uniformly sampled 64 face frames for each video clip. Affect features were extracted by Emonet [33], where the feature included 5-class emotions, arousal, and valence. For audio features, Mel Spectrograms were extracted by OpenSmile toolkit [34]. Raw audio waveforms were also used in our experiments.
Protocols. Inspired by [35, 36], we treated each dataset as a domain. To evaluate the models’ cross-domain generalization capacity and alleviate domain information leakage, all the preprocessed data including original training and test data from each dataset was used for either training or testing. Note that the Box of Lies dataset was only used for testing as many samples were filtered out due to their unsatisfactory quality. The experiments were conducted on the single-to-single domain (e.g., R to B1 stands for training on Real-life Trial (R) and testing on Bag of Lies (B1)) and multi-to-single domain (e.g., R&M to B2 stands for training on Real-life Trial (R) and MU3D (M) and testing on Box of Lies (B2)).
Model Selection. Models for audio and visual modalities were selected to fit the data volume. For face frames, we adopted ResNet18 [31] and Gate Recurrent Unit (GRU) [40] models for facial feature extraction and temporal modeling, respectively. Two-layer multilayer perception (MLP) [41] models were used for AUs, gaze, and affect feature representation. For the audio-based Mel spectrogram, we used the ResNet18 [31] model for time-frequency feature representation. For audio waveforms, the Wave2Vec [42] model was applied for audio feature extraction.
Experimental Setting. Our proposed method was implemented with Pytorch. The ImageNet pretrained models (e.g., ResNet18) for classification were trained on the benchmark datasets using SGD optimizer with the initial learning rate (lr), momentum, and weight decay (wd) were 1e-3, 0.9, and 5e-5, respectively. We trained models with a maximum of 30 epochs and batchsize 32 on a single Nvidia V100 GPU. As for the fusion models (e.g., Atten-Mixer on face frames and Mel Spectrogram), Adam optimizer with initial lr=1e-3 and wd=5e-5 was used. The models were trained with batchsize 16 for a maximum of 30 epochs.
TABLE III: The fusion results of single-to-single cross-domain generalization accuracy (%).
Modality & Inputs Fusion Postion Fusion Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
Score-level Average 53.23 41.58 51.88 64.49 62.38 50.62 57.94 53.23 62.38 55.30
Concat 51.08 55.45 51.88 54.21 58.42 51.25 57.01 50.15 61.39 54.54
SE-Concat 53.85 60.40 51.25 55.14 58.42 50.62 56.07 51.38 65.35 55.83
Cross-Atten 55.38 61.39 52.19 51.40 60.40 51.25 55.14 56.31 60.40 55.98
MLP-Mixer 55.08 48.51 53.44 56.07 58.42 53.75 55.14 59.08 59.41 55.43
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.92 59.41 57.94 63.37 53.75 53.75 60.75 56.00 61.39 58.14
Score-level Average 53.23 49.50 51.88 50.47 59.41 53.75 65.42 56.31 49.50 54.39
Concat 50.77 53.47 50.47 62.38 51.56 51.88 54.21 52.62 58.42 53.98
SE-Concat 50.15 44.55 51.40 61.39 53.12 52.50 65.42 56.31 66.34 55.69
Cross-Atten 54.46 51.49 55.14 58.42 51.25 52.19 63.55 56.62 66.34 55.95
MLP-Mixer 52.31 55.45 57.94 63.37 53.12 51.25 64.49 57.85 62.38 57.57
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.54 55.45 56.07 61.39 50.94 53.12 67.29 57.85 64.36 58.22
Score-level Average 49.85 58.42 54.06 45.79 60.40 50.62 49.53 56.00 63.37 54.23
Concat 49.54 53.47 47.66 61.39 53.44 51.88 57.01 50.15 58.42 53.66
SE-Concat 50.15 48.51 55.14 60.40 53.12 50.00 57.01 49.85 63.37 54.17
Cross-Atten 53.23 44.55 57.94 55.45 54.69 54.06 63.55 51.69 64.36 55.50
MLP-Mixer 49.54 57.43 50.47 63.37 54.06 53.12 59.81 55.08 69.31 56.91
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.52 54.46 57.94 61.39 53.12 51.56 69.16 58.15 64.36 58.18
Score-level Average 52.00 58.42 51.88 53.27 58.42 50.94 57.01 53.54 61.39 55.20
Concat 53.23 58.42 55.14 61.39 52.50 52.19 56.07 54.15 60.40 55.94
SE-Concat 51.38 58.42 51.40 62.38 52.81 50.94 59.81 54.77 52.48 54.93
Cross-Atten 51.08 48.51 55.14 60.40 53.12 53.44 60.75 56.31 60.40 55.46
MLP-Mixer 55.69 46.53 44.86 63.37 51.56 50.94 64.49 56.00 60.40 54.87
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 60.40 57.01 64.36 53.44 51.25 67.29 56.00 62.38 58.58
4.3 Cross-domain Testing with Unimodal Features
In this subsection, we present the benchmark results of cross-domain testing by investigating unimodal features to evaluate their generalization capacities. For clarity, we use “visual (face frames)” to indicate face inputs and “visual (AU/gaze/affect)” to indicate behavior inputs.
Single-to-Single Domain. Specifically, the models were trained on one dataset from one domain and tested on the other dataset from another domain. The experiments were conducted on the four public datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M). As shown in Tabel I, for visual modalities, we extracted the most adopted visual features including face frames, and behavior features such as AUs, gaze and affect. For audio modality, Mel spectrogram and waveform were extracted. We also applied several different backbone networks as audio and visual encoders. We can observe that R and B1 datasets generalized the best on B2, and M generalized the best on R. Note that the B2 dataset was not adopted as a source domain dataset as the original dataset had too much noise and we cleaned it only for testing. On average, we can observe that the best result was achieved by using visual (AU+gaze+affect) features.
Multi-to-Single Domain. Here we fully evaluate the performance of multi-to-single cross-domain generalization using unimodal features. We conducted experiments for different domain sampling strategies. As shown in Table II, for domain-simultaneous training, the best generalization performance is achieved by training on the Bag of Lies and MU3D datasets and testing on the Real-life Trial dataset (69.16%), which was also the case for domain-by-domain training strategy (67.29%). For the domain-alternating strategy, the best result was observed when transferring to the Box of Lies dataset using the rest three datasets for training (65.85%). The results showed that the best generalization performances were obtained when transferring to the real-world dataset and the gameshow dataset by training on the lab-based datasets. This was because the lab-based datasets (Bag of Lies and MU3D) are relatively clean compared to the real-world dataset (Real-life Trial) and the gameshow dataset (Box of Lies). However, in the opposite case, the generalization performance degraded, for example, R&M to B1 and R&B1 to M. Different domain sampling strategies reached their best average performance on different input features and backbone networks. To be specific, for both domain-simultaneous and domain-by-domain strategies, models trained on visual (AU+gaze+affect) features reached their highest accuracies, which were 55.82% and 56.67%, respectively. Using the domain-alternating strategy, the best accuracy of 56.25% was achieved by training on visual (face frames) features. We can observe that models trained on visual modalities outperformed those trained on audio modalities across all the generalization strategies. This may be due to the rich deceptive cues captured by visual modalities in the publicly available datasets.
4.4 Domain-Simultaneous with Gradient Reversal Layer (GRL)
Following the implementation by Ganin et al. [43], we compared the multi-to-single domain generalization accuracies w/w.o GRL. GRL was proposed to mitigate the domain shift issue by manipulating the training gradients. It worked by acting as an identity transform in forward propagation and multiplying the gradient by a certain negative constant during the backpropagation without having trainable parameters. GRL was inserted between encoders and domain classifiers, which was easy to implement. As GRL is a widely adopted method for domain generalization, it is investigated to show its effectiveness for the deception detection task. We selected domain-simultaneous as the baseline and added GRL to the original network with the same training setups. The average accuracies were reported in Fig. 3, where different types of visual and audio features and methods were compared. Training with GRL, the performance of ResNet18 and ResNet18+GRU models using visual (face frames) features and Wave2Vec model using waveform were enhanced. However, we observed that MLP models using visual (AU/gaze/affect) features and the ResNet18 model using Mel spectrograms degraded in performance. Generally, ResNet18 trained with GRL performed better than MLP for visual modality, and Wave2Vec trained with GRL boosted the performance and surpassed the model trained on the Mel spectrogram.
Figure 3: Performance comparisons of Domain-simultaneous training w/ and w/o Gradient Reversal Layer (GRL).
TABLE IV: The fusion results of multi-to-single cross-domain generalization accuracy (%) for different generalization strategies.
Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
Domain-Simultaneous
Score-level Average 53.23 43.56 57.43 50.94 51.40 58.42 48.51 51.93
Concat 52.00 56.44 60.40 51.88 51.40 62.38 57.43 55.99
SE-Concat 55.69 58.42 44.55 52.50 56.07 58.42 58.42 54.87
Cross-Atten 51.08 52.48 55.45 51.56 57.01 60.40 57.43 55.06
MLP-Mixer 53.85 43.56 62.38 52.19 56.07 61.39 59.41 55.55
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.23 57.43 63.37 52.19 57.01 63.37 62.38 58.43
Score-level Average 49.54 55.45 52.48 51.25 54.21 59.41 55.45 53.97
Concat 49.54 54.46 49.50 53.75 44.86 63.37 54.46 52.85
SE-Concat 52.00 58.42 42.57 51.56 54.21 58.42 63.37 54.36
Cross-Atten 50.46 52.48 60.40 53.12 52.34 58.42 59.41 55.23
MLP-Mixer 53.54 57.43 52.48 50.31 52.34 60.40 47.52 53.43
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 64.36 56.44 53.75 58.88 59.41 58.42 58.14
Score-level Average 48.62 58.42 56.44 50.94 51.40 56.44 53.47 53.68
Concat 48.31 53.47 56.44 50.31 56.07 63.37 57.43 55.06
SE-Concat 49.85 50.50 47.52 52.19 49.53 57.43 61.39 52.63
Cross-Atten 51.38 63.37 61.39 51.25 48.60 59.41 58.42 56.26
MLP-Mixer 55.69 59.41 57.43 53.75 50.47 61.39 50.50 55.52
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 55.45 58.42 54.37 52.34 61.39 60.40 56.78
Score-level Average 54.77 57.43 57.43 52.81 60.75 58.42 56.44 56.86
Concat 54.46 54.46 47.52 53.12 53.27 61.39 56.44 54.38
SE-Concat 53.23 55.45 54.46 52.50 59.81 56.44 58.42 55.76
Cross-Atten 48.62 57.43 50.50 53.12 48.60 58.42 56.44 53.30
MLP-Mixer 48.92 48.51 59.41 52.81 54.21 57.43 51.49 53.25
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 52.31 59.41 58.42 52.50 58.88 61.39 60.40 57.62
Domain-Alternating
Score-level Average 55.38 57.43 58.42 50.94 57.94 50.50 61.39 56.00
Concat 55.08 58.42 63.37 51.56 62.62 62.38 53.47 58.13
SE-Concat 50.46 49.50 63.37 52.19 57.01 65.35 56.44 56.33
Cross-Atten 55.08 59.41 60.40 52.50 48.60 58.42 61.39 56.54
MLP-Mixer 57.54 63.37 63.37 50.94 65.42 52.48 56.44 58.51
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.62 61.39 63.37 51.56 58.88 62.38 58.42 58.95
Score-level Average 50.15 58.42 60.40 50.62 51.40 60.40 58.42 55.69
Concat 53.23 62.38 54.46 51.25 59.81 54.46 54.46 55.72
SE-Concat 50.15 52.48 64.36 50.31 59.81 55.45 50.50 54.72
Cross-Atten 52.92 57.43 49.50 52.81 65.42 53.47 66.34 56.84
MLP-Mixer 51.69 58.42 57.43 51.25 62.62 58.42 56.44 56.61
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.23 63.37 56.44 51.25 63.55 59.41 56.44 57.67
Score-level Average 51.38 56.44 54.46 50.00 64.49 61.39 49.50 55.38
Concat 49.85 62.38 58.42 51.88 61.68 55.45 51.49 55.88
SE-Concat 50.46 62.38 58.42 50.00 57.94 59.41 65.35 57.71
Cross-Atten 49.85 62.38 50.50 53.44 58.88 56.44 56.44 55.42
MLP-Mixer 50.15 55.45 58.42 50.31 58.88 61.39 57.43 56.00
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 50.15 59.41 50.50 52.19 58.88 60.40 61.39 56.13
Score-level Average 56.00 52.48 59.41 51.88 51.40 60.40 56.44 55.43
Concat 51.69 58.42 58.42 52.81 57.01 61.39 56.44 56.60
SE-Concat 56.31 59.41 58.42 50.31 53.27 61.39 55.45 56.37
Cross-Atten 49.54 67.33 54.46 51.56 66.36 52.48 60.40 57.45
MLP-Mixer 50.15 63.37 60.40 51.56 62.62 61.39 61.39 58.70
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 66.34 60.40 51.88 59.81 61.39 62.38 59.13
Domain-by-Domain
Score-level Average 54.77 46.53 59.41 50.62 57.94 59.41 58.42 55.30
Concat 53.85 53.47 63.37 51.88 57.94 65.35 58.42 57.75
SE-Concat 56.62 44.55 63.37 52.19 54.21 58.42 59.41 55.54
Cross-Atten 54.15 49.50 62.38 54.69 57.01 60.40 63.37 57.36
MLP-Mixer 57.54 62.38 71.29 51.25 54.21 62.38 62.38 60.20
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 57.54 58.42 63.37 52.19 69.16 60.40 64.36 60.78
Score-level Average 52.00 50.50 58.42 51.56 57.94 61.39 60.40 56.03
Concat 50.46 59.41 62.38 52.19 60.75 56.44 57.43 57.01
SE-Concat 55.38 48.51 58.42 52.81 62.62 57.43 59.41 56.37
Cross-Atten 52.31 53.47 58.42 51.25 64.49 58.42 57.43 56.54
MLP-Mixer 53.23 60.40 62.38 52.19 61.68 58.42 58.42 58.10
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.85 66.34 63.37 53.75 58.88 64.36 63.37 60.56
Score-level Average 48.62 59.41 51.49 51.25 60.75 62.38 63.37 56.75
Concat 57.23 60.40 56.44 51.25 56.07 57.43 56.44 56.47
SE-Concat 55.08 60.40 59.41 54.06 60.75 61.39 57.43 58.36
Cross-Atten 50.15 56.44 61.49 57.50 59.81 64.36 58.42 58.31
MLP-Mixer 52.62 56.44 58.42 52.50 60.75 61.39 58.42 57.22
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 56.44 57.43 54.06 66.36 64.36 60.40 58.68
Score-level Average 53.85 51.49 65.35 53.75 58.88 57.43 58.42 57.02
Concat 52.92 52.48 55.45 53.75 62.62 62.38 56.44 56.58
SE-Concat 51.38 60.40 58.42 51.56 57.01 62.38 59.41 57.22
Cross-Atten 51.80 52.48 62.38 53.12 59.81 64.36 60.40 57.76
MLP-Mixer 53.85 55.45 60.40 50.94 59.81 64.36 60.40 57.89
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 61.39 59.41 55.00 58.88 59.41 60.40 58.60
4.5 Cross-domain Testing with Multimodal Fusion
Here we present multimodal fusion results of cross-domain testing to evaluate models’ generalization capacities.
Single-to-Single Domain with Fusion. In this section, we conducted experiments of single-to-single cross-domain testing. Two types of fusion positions were involved including score-level and feature-level fusion. For feature-level fusion, multiple fusion methods were adopted such as simple concatenation, SE-Concat [44], Cross-Atten [37], MLP-Mixer [19], and Attention-Mixer fusion (Ours). To be specific, simple concatenation refers to the concatenation of the extracted features before the input to the classifier. SE-concat stands for the SE attention applied to the concatenated features. Cross-Atten means the crossmodal attention among input features by using the attention mechanism from the Transformer. MLP-Mixer uses the method in  [19] on the extracted features. The modalities and input include three types, visual (face frames), visual (AU+gaze+affect), and audio (Mel spectrogram). It results in four combinations of either two or three types of inputs. As shown in Table IV, among each type of input combination, the proposed Attention-Mixer fusion (Atten-Mixer) achieved the best results. The rest fusion methods showed comparable results, which were less performed than Atten-Mixer.
Multi-to-Single Domain with Fusion. We benchmarked multi-to-single cross-domain generalization with different fusion methods by three types of cross-testing strategies. In total, seven sub-experiments were conducted for different domain combinations. The results are shown in Tabel IV. On average, the best accuracies were 58.43%, 59.13%, and 60.78% on domain-simultaneous, domain-alternating, and domain-by-domain strategies, respectively. Among these, for both domain-simultaneous and domain-by-domain strategies, the best results were achieved by taking visual (face frames) and visual (AU+gaze+affect) features as input, while the best result for domain-alternating was achieved by using visual (face frames), visual (AU+gaze+affect) and audio (Mel spectrogram) features. Taking a close look at the average fusion results, for each type of modality input, Atten-Mixer achieved the best among the six fusion methods. showing the effectiveness of the proposed method. By using Atten-Mixer, in general, the average result showed slightly better when using the domain-by-domain strategy and taking visual (face frames) and visual (AU+gaze+affect) features as input. To sum up, the results showed that on the current publicly available datasets, visual features were better when it came to multi-to-single cross-domain generalizability. However, the performance differences were not a large gap, and there were also no significant differences by comparing two-to-one domain and three-to-one domain cross-testing performances.
Figure 4: Ablation study for attention-mixer layers. The number of layers 4, 5, 6, and 7 are compared. The modality and inputs for A, B, C, and D are in line with those in Table III from the top to the bottom.
Ablation Study for Attention-Mixer Fusion Module. We conducted an ablation study for the proposed attention-mixer fusion module with the changes in the number of attention-mixer layers. The experiments were conducted on single-to-single domain testing, where the average accuracies were compared. As shown in Fig. 4, the number of attention-mixer layers was set to 4, 5, 6, and 7. The modalities and inputs in Table III are compared, where “A” had the inputs of Visual (Face frame) + Visual(AU+Gaze+Affect), “B” had Visual (Face frames) + Audio(Mel spectrogram), “C” had Visual (AU+Gaze+Affect) + Audio (Mel spectrogram), and “D” had Visual (Face frame) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram). The results showed that models with 6 attention-mixer layers achieved the best average accuracies, followed by 7 attention-mixer layers.
4.6 Discussion
We can observe that the general performance of cross-domain deception detection is unsatisfactory because it is challenging to reduce the domain gap between each dataset. The domain generalization ability of widely-adopted methods was relatively weak using either audio or visual features. Different domain sampling strategies worked well for different audiovisual features. Fusing multiple modalities is able to mitigate the problem. However, the performance still needs to be improved.
Ethical Consideration. Developing deception detection using AI should emphasize respecting privacy, minimizing psychological harm, preventing discrimination, promoting transparency, etc. Researchers should follow appropriate regulations to develop and deploy AI systems for deception detection. Potential misuses and negative impacts include invasion of privacy, discrimination, erosion of trust, etc. Mitigating these risks requires responsible practices from researchers and developers.
5 Conclusion
In this paper, we benchmark the cross-domain generalization performance for deception detection on publicly available datasets. We compare the single-to-single domain and multi-to-single domain generalization performances, where three strategies are used including domain-simultaneous, domain-alternating, and domain-by-domain. We also investigate the effectiveness of the gradient reversal layer for domain-simultaneous strategy. Moreover, we propose the Attention-Mixer fusion method to alleviate the domain shift issue and boost the performance. Future works for deception detection are encouraged to propose better methods to improve the domain generalizability on audio-visual deception detection task.
Acknowledgments.  This work was carried out at the Rapid-Rich Object Search (ROSE) Lab, Nanyang Technological University, Singapore. The research is supported by the DSO National Laboratories, under project agreement No. DSOCL21238.